OpenStudy (anonymous):

if a,b,c are positive and unequal.show that the value of the given determinant is negative.

OpenStudy (anonymous):

I a b c I I b c a I I c a b I

OpenStudy (ash2326):

Let's find the determinant \[\Delta=a(bc-a^2)-b(b^2-ac)+c(ab-c^2) \] Let's simplify it \[abc-a^3-b^3+abc+abc-c^3\] We get \[3abc-a^3-b^3-c^3\]

OpenStudy (ash2326):

Let me think, how we can prove this to be always negative!!

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

ash look -a^3 = always negative -3^3 = -27 so i think its proved :)

OpenStudy (ash2326):

That's correct but we need to prove that \[abc<a^3+b^3+c^3\] for all a's, b's and c's

OpenStudy (anonymous):

take an example |1 2 3| |3 4 1| |4 2 1|

OpenStudy (anonymous):

we need to find its negative value and thats with a prove.

OpenStudy (ash2326):

No, just one example is not enough, we need to prove it for all real numbers

OpenStudy (anonymous):

yah .. @ash2326 is right @annas guys help me plz

OpenStudy (anonymous):

yeah i m trying

OpenStudy (anonymous):

plz ..

OpenStudy (anonymous):

http://www.scribd.com/doc/88441256/19/Applications-of-Determinants-and-Matrices check the example number 30

OpenStudy (anonymous):

we'll have to perform little row and column operations

OpenStudy (anonymous):

taht link is taking too much time ..

OpenStudy (anonymous):

its a bit heavy site

OpenStudy (anonymous):

@shruti if you want i can put a screen shot of the solution ?

OpenStudy (anonymous):

yes sure. i think @ash2326 left.

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

fine

OpenStudy (ash2326):

I'm here, I tried something but it wasn't correct :'(

OpenStudy (anonymous):

:(aww...

OpenStudy (anonymous):

check the attachment

OpenStudy (anonymous):

no worries ash :) you did the best help you are the most awesome helper